OpenStudy (mikezack123):

Given the equation y − 3 = (x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. x − 2y = −12 y = 1/2x y = 1/2x + 6 y = 1/2x + 9

OpenStudy (magbak):

YES I am AM HERE MY PULPI

OpenStudy (mikezack123):

lol. Gandalf, you're here!

OpenStudy (magbak):

Yes but one thing this question has a error.

OpenStudy (mikezack123):

oops, sorry. here you go. Given the equation y − 3 = 1/2(x + 6) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (magbak):

It is B because slop intercept form is y=mx + b m= 1/2 and their is no y intercept.

OpenStudy (phi):

the first step is "distribute" the 1/2 in y − 3 = 1/2(x + 6) can you do that ?

OpenStudy (phi):

if you think of the parens as a package, and you have 1/2 of a package, you have 1/2 of everything in the package.

OpenStudy (phi):

the 2nd step is add +3 to both sides.

OpenStudy (mikezack123):

@phi and then...

OpenStudy (phi):

the first step is "distribute" the 1/2 in y − 3 = 1/2(x + 6) can you do that ?

OpenStudy (phi):

distribute means multiply 1/2 times each thing inside the parens

OpenStudy (mikezack123):

1/2x+ 3

OpenStudy (phi):

multiply 1/2 means write 1/2 * in front of each thing so you get y - 3 = 1/2 x + 3 2nd step is : add +3 to both sides

OpenStudy (phi):

add +3 means write +3 on both sides, then simplify

OpenStudy (mikezack123):

y=1/2x+6

OpenStudy (magbak):

he is correct

OpenStudy (phi):

now go over the steps, and see if they make sense. All these problems are the same, and once you figure it out, you can do all of them

OpenStudy (mikezack123):

Awesome! Thanks @phi !!! :)